OpenStudy (anonymous):

Solve cos (3 theta)= 1

OpenStudy (xapproachesinfinity):

\(cos3\theta=1~ means ~3\theta=2k\pi ~or~ 3\theta=\pi+2k\pi~ ~; k=0,\pm1, \pm2,\pm3..... \)

OpenStudy (xapproachesinfinity):

solve for theta

OpenStudy (anonymous):

Would it be theta= cos^-1(1)

OpenStudy (xapproachesinfinity):

nope that won't quite work and by the way it is 3theta=cos^-1(1) Not theta

OpenStudy (anonymous):

So what do I due then?

OpenStudy (xapproachesinfinity):

oh by the way i made a mistake is should be 3theta=2kpi the other part is wrong! i was thinking something else

OpenStudy (xapproachesinfinity):

you just solve for theta here \(3\theta=2k\pi\)

OpenStudy (anonymous):

What does k stand for?

OpenStudy (xapproachesinfinity):

that's all you can do if you specify an interval we could give some exact values for theta

OpenStudy (xapproachesinfinity):

i wrote it about k is any integer

OpenStudy (xapproachesinfinity):

above*

OpenStudy (anonymous):

I put it in the cal. And got 0

OpenStudy (xapproachesinfinity):

you can't put it calc your answer is \(\large\theta=\frac{2k\pi}{3}\) that's all you can do unless you specify what interval you want like [0, 2pi]

OpenStudy (anonymous):

The intervals are 0 equal or greater than theta equal or greater than 2pi

OpenStudy (xapproachesinfinity):

you see you said just solve, you have to include every detail of the problem without missing any piece so we know where are headed! if you just say solve this! that's mean you are looking for a general solution! now between 0 an 2pi there are only two values that make cos=1 and those are 0 and 2pi themselves in our problem we have 3theta=0 or 3theta=2pi meaning that theta=0 or theta=2pi/3

OpenStudy (anonymous):

That's it!

OpenStudy (xapproachesinfinity):

what else you want that's all. you have two possible answers for theta 0 or 2pi/3

OpenStudy (anonymous):

Thanks

OpenStudy (xapproachesinfinity):

make sure next time to include whatever the problem states! don't miss any thing

OpenStudy (xapproachesinfinity):

you are welcome